OpenStudy (anonymous):

Water is to be pumped to the top of a building that is 366 m high. If the density of water is 1.00 x 103 kg/m3, what amount of pressure is needed in the water line at the base of the building to raise the water to this height

OpenStudy (anonymous):

I'm assuming that you aren't considering any head losses and that this is a straightforward hydrostatics problem. If that's the case, the hydrostatic pressure of any body of liquid is related to its density and its height. P=p*g*z where p = density, g = gravity and z = height

OpenStudy (michele_laino):

hint: here we have to apply the theorem of Bernoulli: \[{z_1} + \frac{{{P_1}}}{{\delta g}} = {z_2} + \frac{{{P_2}}}{{\delta g}}\] where \delta is the water density, and g is the gravity: g=9.81 m/sec^2 or g=32 feet/sec^2